OpenStudy (anonymous):

-tan^2x+sec^2x=1 proving identity help?

hero (hero):

\[-\tan^2x +\sec^2x = 1 \\\sec^2x = 1+\tan^2x \\\frac{1}{\cos^2x} = \frac{\cos^2x}{\cos^2x} + \frac{\sin^2x}{\cos^2x} \\1 = \cos^2x + \sin^2x \\1 = 1 \]